CHW8-Ch10 Item 4 When at a certain rate, the gauge Value Units
Answer #1

cramster-equation-2011111125243634303471

A l vlー/1212

r2 vl 一-1,2 2 rl

cramster-equation-2011111125342634303472

כן2 pl - p2 グ4  

V_2 =

sqrt((31.5Ã10^3 - 25.6Ã10^3)/(0.5Ã1000 (1 - (2.4/3.4)^4)))

= 3.96197 m/s

A2*V2 = pi *( 2.4*10^-2)^2*3.96197 = 0.00716941201 =7.169412*10^-3 m3 /s

let me know in a comment if something goes wrong